# Vision Transformers Need More Than Registers

### v1

@@ 1 Introduction
Vision Transformers (ViTs) [8] have become the de facto standard for image recognition [7]. More importantly, they serve as general-purpose feature extractors across various specific vision tasks [14, 4], functioning as a frozen foundation model pre-trained on large-scale data to embed images into feature representations, enabled by their scalability in data and model size. More broadly, this generic ViT feature extractor can adapt to various supervision methods during pre-training, with different approaches exhibiting characteristics particularly suited to diverse downstream tasks. Specifically, supervised methods—such as training ViTs with fully-supervised classification labels or text-supervised image–text pairs (e.g., in models like CLIP [29])—produce dense features for open-vocabulary tasks, and function as visual encoders for large vision-language models (LVLMs) [23]. Alternatively, self-supervised methods [2, 1, 27], particularly the DINO [1] model trained solely on images, demonstrate the potential for object and part discovery, making them applicable to unsupervised segmentation tasks [33].
@@ 2 Related Work
Artifacts in text-supervised ViTs (CLIP-type models [29]). Recent advances in vision–language contrastive pretraining [29, 19] have enabled CLIP models to produce dense predictions beyond image-level classification. MaskCLIP [51] first showed that CLIP features can yield zero-shot semantic segmentation via pixel–text alignment. However, later studies [38, 18, 47, 12, 20, 40] found that although ViTs surpass ResNets in model capacity and classification accuracy, they perform worse on dense alignment tasks. To mitigate this misalignment, existing works either (1) modify the final attention layers [38, 18, 47, 12, 20, 42] or (2) introduce additional alignment training [40, 3]. In contrast, our method tackles the problem directly during pretraining, avoiding both architectural changes and post-hoc fine-tuning, and fundamentally preventing the emergence of lazy behavior in text-supervised ViTs.
@@ 6.1 Experiment Settings
We first verify the elimination of artifacts in patch score (Sec. 6.2) and validate our proposed method on three training methods: fully supervised (Sec. 6.3), text-supervised (Sec. 6.4), and self-supervised (Sec. 6.5), and examine multiple downstream tasks for ViT under different supervision, including object discovery [33, 1], zero-shot semantic segmentation [29, 34, 43], open-vocabulary object detection [17], instance segmentation [40] and coarse segmentation [1].
@@ 7 Conclusion
Supplementary Material
In the supplementary material, we provide additional information regarding,
  • Implementation Details, Dataset Information and Evaluation Metric (In Section 8).
  • More Comparison (In Section 9).
  • More Qualitative Results about Patch Score (In Section 10).
  • Detailed Analysis about Norm Stratification Phenomenon in DeiT (In Section 11).
  • Detailed Analysis about High-Norm Token (In Section 12).
  • Discussion, Limitation and Future Work (In Section 13).
8 Implementation Details, Dataset Information and Evaluation Metric
In this section, we provide a detailed overview of our implementation details, dataset information, and evaluation metrics utilized in the main experiments. We present the details categorized by different pretraining approaches.
8.1 Text-supervised LaSt-ViT
Implementation Details. Tab. 10 provide our details of training hyper-parameter settings for different ViT variants. We adopt open-sourced OpenCLIP [15] as our code base and train our models on LAION-400M [31]. Given the extensive computational resources required for contrastive learning, we choose to fine-tune models from multiple open-source weights [29, 34, 43].
[…]
Training Config LaSt-ViT-B LaSt-ViT-L
2242 3362
batch size 5376 4k
seen samples 0.64B 0.64B
optimizer AdamW AdamW
base learning rate 5e-4 2e-3
weight decay 0.05 0.02
optimizer momentum $\beta_{1}$ 0.9 0.9
optimizer momentum $\beta_{2}$ 0.98 0.98
learning rate schedule cosine decay cosine decay
warmup steps 1000 2000
warmup schedule linear linear
grad clip norm 1.0 1.0
precision amp bfloat16 amp bfloat16
Table 10: Text-supervised training settings for variants.
8.2 Self-supervised LaSt-ViT
Implementation Details. We strictly follow the DINO [1] framework as our baseline and adopt the same training settings as DINO.
Dataset Information and Evaluation Metric. Following the DINO [1] framework, we train models on ImageNet-1k [7]. Following Register [6], we select object discovery as the benchmark. We employ the Correct Localization (CorLoc) metric, which measures the percentage of correctly localized bounding boxes. A predicted box is deemed correct if its intersection over union (IoU) score exceeds $0.5$ with one of the labeled object bounding boxes. We compare our method on three commonly used object discovery datasets, including PASCAL VOC 2012 [9] (VOC12), PASCAL VOC 2007 [9] (VOC07) and COCO-2012 [22] (COCO).
8.3 Label-supervised LaSt-ViT
Implementation Details. Tab. 11 provide our details of training hyper-parameter settings for different ViT variants. We adopt the official PyTorch [25] implementation of ViT which utilizes DeiT [35]’s training recipe.
Dataset Information and Evaluation Metric. We report classification performance on ImageNet-1K [7] and object detection and instance segmentation performance on the downstream COCO [22] dataset. Additionally, we provide the coarse segmentation on VOC12 and adopt mIoU as evaluation metric. To compute the coarse segmentation, suppose we have already obtained patch score A $\in$ $\mathbb{R}^{H\times W}$ for a given image. We then threshold the patch score map by setting its mean score as the threshold, retaining only patches with a patch score greater than the mean, which are set to 1. The remaining map $\in$ $\{0,1\}^{H\times W}$ forms a segmentation map, which is then compared with all ground truth foreground regions to compute the mIoU.
Training Config LaSt-ViT-S LaSt-ViT-B
2242 2242
batch size 4096 4096
patch size 16 16
epoch 100 100
optimizer AdamW AdamW
base learning rate 1e-3 1e-3
weight decay 0.3 0.3
optimizer momentum $\beta_{1}$ 0.9 0.9
optimizer momentum $\beta_{2}$ 0.95 0.95
learning rate schedule cosine decay cosine decay
warmup steps 500 2000
warmup schedule linear linear
Table 11: Label-supervised training settings for variants.
9 More Comparison
More Comparison on open vocabulary detection. Tab. 12 presents a comparison with previous state-of-the-art methods on the OV-COCO benchmark for open-vocabulary detection. A significant performance gap is observed between ResNet-based and Transformer-based approaches. LaSt-ViT achieves comparable performance on $\text{AP}_{50}^{\text{novel}}$ relative to the best-performing ResNet method [32], demonstrating the effectiveness of LazyStrike during ViT pretraining.
[…]
OV-RCNN [46] RN50 17.5
RegionCLIP [49] RN50 26.8
ViLD [10] RN50 27.6
Detic [52] RN50 27.8
OV-DETR [45] RN50 29.4
VLDet [21] RN50 32.0
BARON [39] RN50 34.0
OC-OVD [30] RN50 36.6
CORA [41] RN50 35.1
EdaDet [32] RN50 37.8
F-VLM [17] RN50 28.0
F-ViT ViT-B/16 17.5
RO-ViT [16] ViT-L/16 33.0
LaSt-ViT ViT-B/16 33.3
LaSt-ViT ViT-L/14 39.1
Table 12: Results on the OV-COCO benchmark.
10 More Qualitative results about Patch Score
In this section, we illustrate more qualitative comparasion of patch score between the ConvNet and Vision Trasnformer in Fig. 10, Fig. 11, and Fig. 12 and provide a in-depth analysis of how network understand an image.
ConvNet focus more on boundary information than ViT. Unlike in ViT, where the patch scores for the entire object are consistently high, ConvNets typically focus on the object’s edges and the most semantically salient regions. As shown in Fig. 10 (rows 1, 3, 4, 5, and 7), the edges of the object are highlighted, while in Fig. 11 (rows 4, 5, 7, and 9), the most semantically salient regions are illuminated. One potential reason could be that convolutional networks are more effective at extracting edge information, while self-attention facilitates the fusion of features between different parts of an object.
DINO-like self-sup. helps ViT to focus on single object. As shown in Fig. 10 (rows 1, 4, and 8), the main difference between label-supervised LaSt-ViT and self-supervised LaSt-ViT lies in their tendency to gathering semantics. Label-supervised LaSt-ViT tends to capture global semantics from all foreground objects. For example, in Fig. 10 (row 1), the boxes, and in Fig. 10 (row 8), the trees, exhibit uniformly high patch scores. In contrast, self-supervised LaSt-ViT focuses on objects with more prominent and distinct instance features, prioritizing their semantic representation.
LazyStrike effectively and consistently mitigates spatial inconsistency across various supervised ViTs. It eliminates artifacts in the feature map and enables interpretable dense features for both supervised (right) and self-supervised (left) ViTs.
11 Analysis about Norm Stratification
Figure 8: Norm stratification phenomenon in DeiT [35]
Fig 8 illustrates the difference in feature norm between w/ and w/o LazyStrike. LazyStrike effectively eliminates high-norm outliers, as tokens no longer need to diffuse foreground semantics. We also observe an interesting norm stratification phenomenon in DeiT [35]. In this section, we provide a detailed analysis of the norm stratification observed in the DeiT model. Norm stratification refers to the hierarchical organization of token norms within the model, where certain tokens exhibit consistently higher norms than others. To better understand the stratification, we visualize the token norms for the original DeiT model and DeiT with LazyStrike in Fig. 13 and Fig. 14, respectively. These visualizations highlight the differences in how token norms are distributed from last layer and the impact of LazyStrike on stratification.
Figure 9: Patch Score v.s. Feature Norm. For each triplet, we present the original image, patch score, and feature norm. Feature norm and patch score exhibit a correlation, where regions with higher feature norms typically have higher patch scores. Experiments demonstrate that patch score remains disordered from the very beginning of training and appears across models of all sizes, whereas high-norm tokens emerge only in the mid-to-late stages of training in larger models. Simultaneously eliminating artifacts in patch score also removes high-norm tokens, leading us to conclude that high-norm tokens are a specific manifestation of patch score artifacts in larger models during the training.
LaSt-ViT learns to distinguish foreground from background by leveraging feature norms. Specifically, in LaSt-ViT’s feature norm, there is a consistently significant difference between foreground and background norms. As shown in Fig. 13, the left column highlights cases where the background norm is larger, while the right column displays cases where the foreground norm is larger. Unlike the original DeiT, where the feature norm often exhibits meaningless high-norm outliers, LaSt-ViT’s feature norm demonstrates that the network distinguishes foreground and background in a more structured and meaningful manner. This behavior highlights LaSt-ViT’s ability to suppress redundant or irrelevant tokens, enabling it to focus on semantically significant regions of the image and improving overall interpretability and task-specific performance.
High norm does not always correspond to more meaningful semantics. The Register [6] suggests that feature norm represents the extent of global information a token possesses, which holds true within the scope of the register’s discussion. However, we have observed that this is not necessarily the case. After addressing the lazy behavior of ViT, there is no absolute correlation between feature norm and global information. In other words, a low norm can also represent global semantics.
12 Detailed Analysis about High-Norm Token.
High-norm tokens are a distinct manifestation of the lazy behavior of large ViTs during the mid-to-late stages of training. Fig.˜9 simultaneously presents patch scores and feature norms. Feature norm and patch score are correlated, with higher feature norms typically corresponding to higher patch scores. Our experiments show that patch scores remain disordered from the start of training across all model sizes, while high-norm tokens emerge only in the mid-to-late stages of larger models. Eliminating patch score artifacts also removes high-norm tokens, suggesting they are a manifestation of patch score artifacts in later training stages of larger models. A reasonable hypothesis is that high-norm tokens result from accumulated patch score artifacts, leading to uneven gradient updates that reinforce certain activations, especially in larger models during later training stages.
Why DINO-v1 [1] is an exception. Register [6] identified DINO-V1 [1] as an exception, as it did not exhibit the high-norm phenomenon, but did not explain why. Through our deeper understanding of the underlying causes of the high-norm phenomenon, we attribute this to DINO’s local-global self-supervised loss. DINO’s supervision signal comes from local crops, and since we have identified the absence of local signals as a key factor in the emergence of high-norm tokens, the introduction of random local supervision through cropping could help mitigate the underlying cause.
13 Discussion, Limitation and Future Work
Discussion of the artifacts in other model. We have also observed the artifact phenomenon in other sequence modeling models, such as Mamba [37] and decoder-only LLMs [23]. We leave the investigation of other model as future work.
   input    ConvNet [13]    Vision Transformer [1, 8]    Transformer w/ LazyStrike (Ours)
Figure 10: More qualitative results for the comparasion between ResNet and ViT under label-supervision (right) and self-supervision (left). We examine the similarity between CLS token and patch features. Input resolution 880 $\times$ 880.
   input    ConvNet [13]    Vision Transformer [1, 8]    Transformer w/ LazyStrike (Ours)
Figure 11: More qualitative results for the comparasion between ResNet and ViT under label-supervision (right) and self-supervision (left). We examine the similarity between CLS token and patch features. Input resolution 880 $\times$ 880.
   input    ConvNet [13]    Vision Transformer [1, 8]    Transformer w/ LazyStrike (Ours)
Figure 12: More qualitative results for the comparasion between ResNet and ViT under label-supervision (right) and self-supervision (left). We examine the similarity between CLS token and patch features. Input resolution 880 $\times$ 880.
   input    DeiT [35]    DeiT w/ LazyStrike input DeiT [35]    DeiT w/ LazyStrike
Figure 13: More qualitative results for the feature norm between DeiT and DeiT with LazyStrike.
   input    DeiT [35]    DeiT w/ LazyStrike input DeiT [35]    DeiT w/ LazyStrike
Figure 14: More qualitative results for the feature norm between DeiT and DeiT with LazyStrike.

### v2

@@ 1 Introduction
Vision Transformers (ViTs) [7] have become the de facto standard for image recognition [6, 22, 23, 24, 40, 13]. More importantly, they serve as general-purpose feature extractors across various specific vision tasks [12, 4, 9, 45, 33], functioning as a frozen foundation model pre-trained on large-scale data to embed images into feature representations, enabled by their scalability in data and model size. More broadly, this generic ViT feature extractor can adapt to various supervision methods during pre-training, with different approaches exhibiting characteristics particularly suited to diverse downstream tasks. Specifically, supervised methods—such as training ViTs with fully-supervised classification labels or text-supervised image–text pairs (e.g., in models like CLIP [27])—produce dense features for open-vocabulary tasks, and function as visual encoders for large vision-language models (LVLMs) [20]. Alternatively, self-supervised methods [2, 1, 25], particularly the DINO [1] model trained solely on images, demonstrate the potential for object and part discovery, making them applicable to unsupervised segmentation tasks [31].
@@ 2 Related Work
Artifacts in text-supervised ViTs (CLIP-type models [27]). Recent advances in vision–language contrastive pretraining [27, 18] have enabled CLIP models to produce dense predictions beyond image-level classification. MaskCLIP [44] first showed that CLIP features can yield zero-shot semantic segmentation via pixel–text alignment. However, later studies [36, 17, 42, 10, 19, 37] found that although ViTs surpass ResNets in model capacity and classification accuracy, they perform worse on dense alignment tasks. To mitigate this misalignment, existing works either (1) modify the final attention layers [36, 17, 42, 10, 19, 38] or (2) introduce additional alignment training [37, 3] or (3) employ test-time activation editing, such as dynamically shifting high-norm outlier activations into untrained register tokens during inference [15]. In contrast, our method tackles the problem directly during pretraining, avoiding architectural changes, post-hoc fine-tuning, and test-time interventions, fundamentally preventing the emergence of lazy behavior in text-supervised ViTs.
@@ 6.1 Experiment Settings
We first verify the elimination of artifacts in patch score (Sec. 6.2) and validate our proposed method on three training methods: fully supervised (Sec. 6.3), text-supervised (Sec. 6.4), and self-supervised (Sec. 6.5), and examine multiple downstream tasks for ViT under different supervision, including object discovery [31, 1], zero-shot semantic segmentation [27, 32, 39], open-vocabulary object detection [16, 30], instance segmentation [37, 29] and coarse segmentation [1].
@@ 7 Conclusion
8 Acknowledgments
This work is supported in part by the Hong Kong Research Grants Council under Collaborative Research Fund (Project No. HKUC7004-22G) and the National Natural Science Foundation of China under Grant No.62576365.